1. A 30.50 g sample of metal is heated, then placed in a calorimeter containing 65.0 g of water. Temperature of water increases from 20.00 °C to 22.50 °C. The initial temperature of the zinc metal sample was 79.7°C? What is the heat capacity of the metal?
1
Expert's answer
2021-10-05T23:58:28-0400

"Q=mC \Delta T"

Q water + Q metal = 0

Q water = - Q metal

65 × 4.184 × ( 22.50 -20.0 ) =30.50 × C × ( 22.50- 20.0)

679.9 = 1744.6C

C = 679.9/ 1744.6


C=0.3897 J/K